Find dy/dx by implicit differentiation. xey = x - y y = Need Help? Read It Talk to a Tutor
Answer #1

Given that xe = x-y xe” –x+y= 0 Suppose f (x,y) = xe – x+y (xe” –x+y) =e) - 1 ах (xe” –x+y = xe” +1 II a Since the function